OpenStudy (clara1223):

find the limit as x approaches 1 of ((6x^4)-6)/(2x-2)

OpenStudy (freckles):

factoring time!

OpenStudy (clara1223):

\[\lim_{x \rightarrow 1}\frac{ 6x ^{4}-6 }{ 2x-2 }\]

OpenStudy (freckles):

\[6x^4-6 \\ 6(x^4-1) \\ 6(x^2-1)(x^2+1) \text{ by differece of squares }\] can you use difference of squares again bet you will see a common factor on top and bottom

random231 (random231):

yup or if you know the l'hospitals theorem then you could use that too.

OpenStudy (jdoe0001):

keep in mind that \(1 = 1^2, 1^3,1^4, 1^5, 1^1000....\)

OpenStudy (jdoe0001):

hmm \(1^{1000}\) rather... anyhow \((x^4-1) = (x^4-1^4)\) thus

OpenStudy (dinamix):

@clara1223 u must find it =12 this limit